2 bodies are projected vertically upwards from one point with the same initial velocity vo. The second body is projected to seconds after the first.How long after will the bodies meet?

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Chilukuri

let at hight h and after t second from the projection of second body both bodies meet

since first body projected t0 second earlier so

h=v(t+t0) -1/2 *g *(t+t0)2

 

for second body

h=vt-1/2 *g *t2

equate both

 v(t+t0) -1/2 *g *(t+t0)2  =   vt-1/2 *g *t2

simplify

2vt0/g =(2t+t0)t0

t={2v/g  -t0}/2


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin